# spacing=nonfrench

### v2

@@ 6.1.2 Baselines
Model choice rationale: We deliberately use small, minimally-capable models (0.5b–3b parameters) rather than frontier models. This design choice strengthens our thesis: if the coordination mechanism can extract effective performance from weak models, the mechanism itself is valuable—independent of model capability. Using identical model chains across all strategies isolates coordination effects from model effects.
@@ 6.2  Main Results
Conversation provides intermediate performance: The AutoGen-style conversation baseline achieves 12.6% overall, significantly better than hierarchical ($p<0.001$) but far below pressure-field. Notably, conversation solves only easy problems (37.8% on easy, 0% on medium and hard).
@@ 6.3.1 Effect of Temporal Decay
Decay proves essential—without it, solve rate drops:
@@
Table 4: Decay ablation on easy scheduling problems (30 trials each). Disabling decay reduces solve rate by 10 percentage points.
The effect is meaningful: disabling decay reduces solve rate from 96.7% to 86.7%. Without decay, fitness saturates after initial patches—regions that received early patches retain high fitness indefinitely, making them appear “stable” even when they still contain unscheduled meetings. Since greedy selection prioritizes high-pressure regions, these prematurely-stabilized regions are never reconsidered. This validates the Basin Separation result (Theorem 3): without decay, agents remain trapped in the first stable basin they reach, even if lower-pressure basins exist elsewhere in the landscape.
@@ 6.3.2 Effect of Inhibition and Examples
Decay dominates: configurations with decay achieve higher solve rates. Inhibition shows no detectable effect in this domain, possibly because the 50-tick budget provides sufficient exploration without explicit cooldowns.
Figure 2: Ablation study results. Left: feature matrix showing which components are enabled. Right: solve rates for each configuration. Decay provides the largest contribution (+10%), followed by examples (+6.7%). Inhibition shows no detectable effect.
@@ 6.4  Scaling Experiments
6.5  Band and Model Escalation
@@
Model escalation addresses a different failure mode: when exploration within a model’s capability envelope fails to discover productive paths, the system recruits more capable agents. This mirrors the immune system’s escalation from innate to adaptive response: macrophages and neutrophils handle routine pathogens quickly and cheaply, but when these first responders fail to clear an infection, the system escalates to T-cells and antibody-producing B-cells—slower to mobilize but capable of addressing threats invisible to simpler mechanisms. Similarly, model escalation (0.5b $\to$ 1.5b $\to$ 3b) reserves greater reasoning capacity for regions that resist simpler approaches. Each model upgrade resets to exploitation band, giving the more capable model opportunity to exploit solutions invisible to its predecessor before resorting to exploration.
This two-level mechanism—behavioral adaptation within agents (band escalation) and capability escalation across agents (model escalation)—maintains the stigmergic principle: coordination emerges from environment signals (pressure gradients) rather than explicit planning. The system does not “decide” to explore or escalate; it reacts to pressure stagnation, just as ants react to pheromone decay.
@@ 6.6  Difficulty Scaling
    The gap widens with difficulty: On easy problems, pressure-field leads by 48.9 percentage points over conversation (86.7% vs 37.8%). On medium and hard problems, the gap becomes absolute—pressure-field solves problems that no baseline can solve.
@@
    Effect sizes are large: Cohen’s $h$ on easy problems: $h=1.07$ vs conversation, $h=1.97$ vs hierarchical, $h=2.18$ vs sequential/random.
Figure 4: Effect sizes (Cohen’s $h$) for pressure-field versus each baseline on easy problems. The dashed line indicates the “large effect” threshold ($h=0.8$). All comparisons exceed this threshold, with effects ranging from $h=1.07$ (vs conversation) to $h=2.18$ (vs sequential/random).
@@ 6.7  Convergence Speed
Figure 5: Mean ticks to solution on easy problems (solved cases only). Error bars show standard error. Pressure-field converges fastest (17.8 ticks), followed by conversation (29.9 ticks) and hierarchical (40.0 ticks).
@@ 8  Conclusion
Our theoretical analysis establishes convergence guarantees under pressure alignment conditions, with coordination overhead independent of agent count. Empirically, on meeting room scheduling across 1350 trials, we find:
@@
    Pressure-field outperforms all baselines (48.5% vs 12.6% for conversation, 1.5% for hierarchical, all $p<0.001$). Implicit coordination through shared pressure gradients exceeds explicit hierarchical coordination.
@@
    Temporal decay is essential. Disabling it reduces solve rate by 10 percentage points, trapping agents in local minima.
@@ A.4  Baseline Implementation Details
Conversation baseline implementation The conversation baseline implements AutoGen-style multi-agent dialogue with three roles: a Coordinator that selects which region to address, a Proposer that suggests schedule changes, and a Validator that critiques proposals. Each tick proceeds as follows: (1) the Coordinator identifies a high-pressure region (1 Large Language Model call), (2) the Proposer and Validator engage in up to 5 dialogue turns (2 calls per turn), (3) if the Validator approves, the patch is applied. This yields 4–12 Large Language Model calls per tick for a single region, compared to pressure-field’s $K$ parallel calls across $K$ regions. The conversation terminates when the Validator approves or the turn limit is reached. We use fixed sampling parameters (T=0.3, top-p=0.9) for structured dialogue, lower than pressure-field’s exploration bands.
@@ A.7  Metrics Collected
  • token_usage: Prompt and completion tokens consumed (pressure-field only; conversation baseline uses a different LLM client path that does not collect token statistics, precluding cross-strategy comparison)

### v3

@@ 6.1.2 Baselines
Model choice rationale: We deliberately use small, minimally-capable models (0.5b–3b parameters) rather than frontier models. This design choice strengthens our thesis: if the coordination mechanism can extract effective performance from weak models, the mechanism itself is valuable—independent of model capability. Using identical model chains across all strategies isolates coordination effects from model effects. We hypothesize that frontier models (e.g., GPT-4, Claude) would raise absolute solve rates across all strategies while preserving relative rankings: the coordination advantage is orthogonal to model capability, so pressure-field’s $4\times$ improvement over conversation baselines should persist even as the baseline rises.
@@ 6.1.3 Metrics
  • Token efficiency: Total prompt and completion tokens consumed per trial and per successful solve
@@ 6.2 Main Results
Conversation provides intermediate performance: The AutoGen-style conversation baseline achieves 11.1% overall, significantly better than hierarchical ($p<0.001$) but far below pressure-field. Notably, conversation solves only easy problems (33.3% on easy, 0% on medium and hard).
@@ 6.3.1 Effect of Temporal Decay
Decay appears beneficial, though the effect does not reach statistical significance with our sample size:
@@
Table 4: Decay ablation on easy scheduling problems (30 trials each). Fisher’s exact test: $p=0.35$, indicating the 10 percentage point difference is not statistically significant at $\alpha=0.05$.
The observed effect—a 10 percentage point reduction when decay is disabled—is directionally consistent with our theoretical predictions. The non-significant $p$-value ($p=0.35$) reflects both limited sample size and a ceiling effect: the high baseline solve rate on easy problems (96.7% with decay) leaves limited statistical room to detect improvement. The overlapping confidence intervals (83.3%–99.4% vs 70.3%–94.7%) reflect this uncertainty. Without decay, fitness saturates after initial patches—regions that received early patches retain high fitness indefinitely, making them appear “stable” even when they still contain unscheduled meetings. Since greedy selection prioritizes high-pressure regions, these prematurely-stabilized regions are never reconsidered. This mechanism is consistent with the Basin Separation result (Theorem 5.3): without decay, agents may remain trapped in the first stable basin they reach. Larger-scale ablation studies would be needed to establish the statistical significance of decay’s contribution.
@@ 6.3.2 Effect of Inhibition and Examples
Decay shows the largest effect: configurations with decay achieve higher solve rates, though the differences do not reach statistical significance at $n=30$. Inhibition shows no detectable effect in this domain, possibly because the 50-tick budget provides sufficient exploration without explicit cooldowns.
Figure 2: Ablation study results. Left: feature matrix showing which components are enabled. Right: solve rates for each configuration. Decay shows the largest observed effect (+10%), followed by examples (+6.7%), though neither reaches statistical significance at $n=30$. Inhibition shows no detectable effect.
@@ 6.4 Scaling Experiments
6.5 Band and Model Escalation: FM-MAS Symbiosis in Practice
The escalation mechanism demonstrates the FM-MAS symbiosis that pressure-field coordination enables. Traditional approaches face a dilemma: use a large, capable model everywhere (expensive) or use a small, efficient model everywhere (limited). Pressure-field coordination dissolves this dilemma through gradient-driven capability invocation: the coordination mechanism itself determines when greater Foundation Model capability is needed.
@@ 6.5 Band and Model Escalation: FM-MAS Symbiosis in Practice
Model escalation addresses a different failure mode: when exploration within a model’s capability envelope fails to discover productive paths, the system recruits more capable Foundation Models. This is where FM-MAS symbiosis becomes concrete: the Multi-Agent System coordination mechanism (pressure-field) adaptively invokes higher-capability Foundation Models based on pressure signals alone. No explicit task decomposition is needed—the pressure gradient indicates that current capabilities are insufficient, triggering escalation. Model escalation (0.5b $\to$ 1.5b $\to$ 3b) reserves greater reasoning capacity for regions that resist simpler approaches. Each model upgrade resets to exploitation band, giving the more capable model opportunity to exploit solutions invisible to its predecessor before resorting to exploration.
This architecture instantiates the bidirectional relationship between Foundation Models and Multi-Agent Systems:
  • FM contribution: Each model tier provides broad solution coverage without requiring explicit action enumeration. The 3b model can propose patches invisible to the 0.5b model, but both operate through the same interface—observe pressure, propose patch.
  • MAS contribution: The pressure gradient provides the objective criterion for when to escalate. No heuristics about “problem difficulty” are needed; stagnant pressure velocity is sufficient signal. The coordination mechanism manages heterogeneous Foundation Model capabilities without explicit capability models.
This two-level mechanism—behavioral adaptation within agents (band escalation) and capability escalation across agents (model escalation)—maintains the stigmergic principle: coordination emerges from environment signals (pressure gradients) rather than explicit planning. The system does not “decide” to explore or escalate; it reacts to pressure stagnation, just as ants react to pheromone decay. The result is adaptive Foundation Model utilization: cheap models handle easy regions, expensive models are reserved for regions where cheap models stall.
@@ 6.6 Difficulty Scaling
    The gap widens with difficulty: On easy problems, pressure-field leads by 53.4 percentage points over conversation (86.7% vs 33.3%). On medium and hard problems, the gap becomes absolute—pressure-field solves problems that no baseline can solve.
@@
    Effect sizes are large: All pairwise comparisons on easy problems exceed Cohen’s “large effect” threshold ($h>0.8$); see Figure 4 for the full breakdown.
Figure 4: Effect sizes (Cohen’s $h$) for pressure-field versus each baseline on easy problems. The dashed line indicates the “large effect” threshold ($h=0.8$). All comparisons exceed this threshold, with effects ranging from $h=1.16$ (vs conversation) to $h=2.18$ (vs sequential/random).
@@ 6.7 Convergence Speed
Figure 5: Mean ticks to solution on easy problems (solved cases only). Error bars show standard error. Pressure-field converges fastest (17.8 ticks), followed by conversation (29.4 ticks) and hierarchical (40.0 ticks).
@@ 6.8 Final Pressure Analysis
[…]
6.9 Token Efficiency
A natural question is whether pressure-field’s superior solve rate comes at the cost of increased Large Language Model usage. We tracked token consumption across all strategies:
Strategy Prompt Completion Total
Pressure-field 532,106 85,165 617,271
Conversation 154,845 6,626 161,472
Hierarchical 32,662 5,424 38,087
Sequential 40,667 5,239 45,906
Random 41,227 5,253 46,480
Table 11: Average token usage per trial. Pressure-field uses more tokens per trial due to parallel multi-agent execution.
[…]
Strategy Total Tokens Solves Tokens/Solve
Pressure-field 166.7M 131 1.27M
Conversation 43.6M 30 1.45M
Table 12: Token efficiency per successful solve (270 trials each). Despite higher per-trial cost, pressure-field achieves lower cost per solve due to its superior success rate.
Pressure-field requires 1.27M tokens per solve versus conversation’s 1.45M—12% more efficient per successful outcome. The apparent cost disadvantage inverts when accounting for success rate.
Token usage also varies by outcome:
Strategy Solved Trials Unsolved Trials
Pressure-field 318K (n=131) 900K (n=139)
Conversation 78K (n=30) 172K (n=240)
Table 13: Average tokens by outcome. Both strategies use fewer tokens on solved trials (early termination), but pressure-field’s unsolved trials are more expensive due to sustained parallel exploration.
[…]
@@ 7.3 Limitations
However, meeting scheduling is representative of a broader class of resource allocation problems with soft constraints: cloud compute scheduling (VMs to hosts with affinity/anti-affinity rules), logistics optimization (packages to routes with capacity and time windows), and workforce rostering (employees to shifts with skill and availability constraints). These domains share the structural properties that enable pressure-field coordination: (1) decomposable quality metrics (utilization, constraint violations, balance), (2) locality (fixing one allocation rarely breaks distant allocations), and (3) multiple valid solutions (no single optimum, many acceptable configurations). We selected meeting scheduling as the evaluation domain because it is intuitive, reproducible, and exhibits these properties clearly. The coordination mechanism itself—pressure gradients, temporal decay, parallel validation—is domain-agnostic; only the pressure function requires domain-specific design.
@@
7.3.1 Hallucination Filtering and Emergent Trajectories
The validation phase (Phase 2b) serves as a filter for individual Foundation Model hallucinations: patches that increase pressure or violate syntactic constraints are rejected before application. This provides strong guarantees at the individual-action level—no single hallucination can directly harm artifact quality.
However, system-level behavior emerges from sequences of validated patches, and trajectory-level risks remain. Consider a hypothetical scenario: an agent proposes a patch that reduces pressure in region $R_{1}$ by making a change that creates a subtle dependency on region $R_{2}$. The patch is validated (pressure decreased), but subsequent patches to $R_{2}$ now have unpredictable effects on $R_{1}$. The individual patches are all pressure-reducing, but the emergent trajectory creates fragile coupling that was invisible to local validation.
More generally, validation filters based on pressure reduction cannot detect:
  • Coherence drift: Individual improvements that collectively shift the artifact toward an inconsistent state
  • Emergent gaming: Patches that exploit pressure function weaknesses in ways only apparent over multiple steps
  • Dependency accumulation: Gradual introduction of hidden couplings that reduce future improvability
Mitigating trajectory-level risks requires mechanisms beyond local validation: periodic global coherence checks, trajectory logging for post-hoc analysis, and pressure functions that penalize not just local quality but also coupling metrics. These remain areas for future work.
7.3.2 Decay Miscalibration Failure Modes
Temporal decay is critical for escaping local optima (Theorem 5.3), but miscalibrated decay rates introduce distinct failure modes:
Too-fast decay prevents stability. When fitness decays faster than agents can reinforce successful regions, the system cannot maintain any stable configuration. Solved regions immediately become high-pressure again, triggering unnecessary rework. In the limit, excessively fast decay produces perpetual oscillation: agents patch, decay erases, agents re-patch, indefinitely.
Too-slow decay traps in local minima. When fitness decays slower than the exploration timescale, agents cannot escape suboptimal basins. Early patches establish fitness peaks that persist indefinitely, preventing reconsideration even when better solutions exist. This is the failure mode our ablation study suggests: without decay, we observed a 10 percentage point reduction in solve rate, consistent with agents remaining trapped in initial basins (though the effect did not reach statistical significance at $n=30$).
The optimal decay rate depends on problem characteristics: harder problems with deeper local minima require faster decay to enable escape, while problems with fragile solutions require slower decay to maintain stability. Our experiments use fixed decay rates ($\lambda_{f}=0.1$, fitness half-life 5 seconds), which may be suboptimal for some problem instances. Adaptive decay—adjusting rates based on pressure velocity—is a promising direction for future work.
@@ 7.6 Future Work
7.7 FM-MAS Reciprocity: Bidirectional Problem Solving
The intersection of Foundation Models and Multi-Agent Systems is not merely additive—each paradigm solves fundamental problems that the other cannot address alone. This reciprocity suggests that pressure-field coordination represents a genuine synthesis rather than a simple combination.
7.7.1 Foundation Models Solve a MAS Coordination Problem
Traditional Multi-Agent System coordination requires explicit action space enumeration. Generalized Partial Global Planning coordinates tasks through declared capabilities and commitments; SharedPlans reasons about action sequences and their preconditions; organizational models assign roles with specified behaviors. All these approaches assume that designers can enumerate what agents can do and under what conditions.
For open-ended artifact refinement—improving code quality, refining documents, optimizing configurations—this enumeration is intractable. The space of possible improvements is unbounded; no finite action language captures all valid patches. Traditional Multi-Agent System approaches require either (a) restricting the problem to a tractable subset with enumerable actions, or (b) developing domain-specific action representations for each artifact type.
Foundation Models eliminate this enumeration requirement. Their broad pretraining provides implicit coverage of improvement strategies across diverse domains without explicit action specification. An Large Language Model actor can propose patches for code, natural language, structured data, or configuration files using the same interface: observe local context, receive pressure feedback, generate improvement proposals. This “universal actor” capability enables stigmergic coordination on problem classes where traditional Multi-Agent System approaches would require extensive domain engineering.
7.7.2 Multi-Agent Systems Solve a FM Problem
Conversely, Foundation Models face a fundamental problem that Multi-Agent System coordination solves: how to combine multiple outputs coherently. A single Foundation Model produces one response per query; scaling to complex artifacts requires orchestrating multiple generations. Current approaches use ad-hoc combination strategies: voting, ranking, chain-of-thought aggregation, or human-in-the-loop selection.
Pressure-field coordination provides a principled framework for combining Foundation Model outputs. Rather than voting on which response is “best” or ranking outputs by heuristic scores, the pressure gradient defines an objective criterion: accept patches that reduce pressure, reject those that do not. Multiple Foundation Model outputs compete not through popularity or arbitrary ranking, but through their effect on a well-defined quality function.
This framing clarifies why pressure-field coordination outperforms conversation baselines. AutoGen-style multi-agent dialogue is an ad-hoc combination strategy—agents exchange messages and reach conclusions through emergent consensus. Pressure-field coordination replaces emergent consensus with objective gradients: no negotiation is needed when the artifact state adjudicates quality.
7.7.3 Implications for FM-MAS Integration
This reciprocity suggests design principles for future FM-MAS systems:
  1. 1.
    Leverage FM coverage, constrain via MAS gradients. Use Foundation Models for their broad solution coverage, but use Multi-Agent System coordination mechanisms to filter and combine outputs objectively.
  2. 2.
    Prefer stigmergic over explicit coordination. When Foundation Models serve as actors, stigmergic coordination avoids the overhead of explicit protocols that FMs may not reliably follow.
  3. 3.
    Design pressure functions as FM-MAS interfaces. The pressure function is the critical interface: it must capture human intent precisely enough that Foundation Model outputs reducing pressure are genuinely improvements.
Our experimental results—pressure-field achieving $4\times$ higher solve rates than conversation baselines—provide empirical support for these principles. The coordination overhead of explicit dialogue exceeds its organizational benefit when objective gradients are available.
@@ 8 Conclusion
Our theoretical analysis establishes convergence guarantees under pressure alignment conditions, with coordination overhead independent of agent count. Empirically, on meeting room scheduling across 1350 total trials (270 per strategy), we find:
@@
    Implicit coordination outperforms explicit coordination. Pressure-field achieves 48.5% aggregate solve rate—nearly half of all problems solved through local pressure-following alone, with no coordinator, no message passing, and no explicit task delegation. The gap versus hierarchical control ($30\times$) and conversation-based coordination ($4\times$) is both large and highly significant (all $p<0.001$).
@@
    Temporal decay shows beneficial effects. Ablation studies observe a 10 percentage point improvement with decay enabled, consistent with theoretical predictions about escaping local minima, though larger samples would be needed to establish statistical significance.
@@
[…]
Code Availability
[…]
Acknowledgments
The author thanks the anonymous reviewers for their constructive feedback.
Declarations
[…]
Funding
No funding was received for conducting this study.
[…]
Competing Interests
The author has no relevant financial or non-financial interests to disclose.
[…]
Data Availability
All experimental data and analysis scripts are available in the code repository.
[…]
Use of AI Tools
The author used Claude (Anthropic) to provide feedback on draft versions of this manuscript. All analytical decisions, experimental design, and final content were made by the author, who takes full responsibility for the work.
@@ A.4 Baseline Implementation Details
Conversation baseline implementation The conversation baseline implements AutoGen-style multi-agent dialogue with three roles: a Coordinator that selects which region to address, a Proposer that suggests schedule changes, and a Validator that critiques proposals. Each tick proceeds as follows: (1) the Coordinator identifies a high-pressure region (1 Large Language Model call), (2) the Proposer and Validator engage in up to 5 dialogue turns (2 calls per turn), (3) if the Validator approves, the patch is applied. This yields 4–12 Large Language Model calls per tick for a single region, compared to pressure-field’s $K$ parallel calls across $K$ regions. The conversation terminates when the Validator approves or the turn limit is reached. We use fixed sampling parameters (T=0.3, top-p=0.9) for structured dialogue, lower than pressure-field’s exploration bands. Note that this represents one instantiation of conversation-based coordination inspired by AutoGen’s design principles; actual AutoGen deployments may use different role configurations, turn protocols, or termination conditions.
Fairness guarantees for baseline comparison: To ensure the conversation baseline had equal opportunity to succeed:
  • Same models: All strategies use identical model chains (qwen2.5:0.5b $\to$ 1.5b $\to$ 3b) with the same escalation triggers
  • Same problem access: All strategies observe identical schedule state, room capacities, meeting requirements, and constraint information
  • Retry logic: The conversation baseline includes up to 5 dialogue turns per region, allowing the Proposer to refine proposals based on Validator feedback—this is more opportunity for refinement than pressure-field’s single-shot proposals
  • Same tick budget: All strategies receive the same 50-tick limit to solve each problem
  • Output parsing: All strategies use identical response parsing and schedule extraction logic
The only intentional differences are the coordination mechanisms themselves: how regions are selected, how proposals are generated, and how patches are validated.
@@ A.7 Metrics Collected
  • token_usage: Prompt and completion tokens consumed per trial
@@ A.9 Estimated Runtime
Appendix B Pressure Alignment Verification
This appendix verifies that the meeting scheduling domain satisfies the alignment condition (Definition 2) required for convergence (Theorem 1).
B.1 Per-Region Pressure Components
The per-region pressure function includes three components:
  • gap_ratio: Fraction of empty slots in this time block. Strictly local—depends only on meetings scheduled within this region.
  • overlap_count: Number of attendee double-bookings within this time block. Strictly local—the overlap sensor counts attendees with multiple meetings in this block only, not across blocks.
  • utilization_variance: Variance in room utilization within this time block. Strictly local—measures balance across rooms for meetings in this region.
The unscheduled_count component is added to total pressure only, not to per-region pressure. This design choice ensures separability.
B.2 Coupling Analysis
For the alignment condition, we need $|P_{j}(s^{\prime})-P_{j}(s)|\leq\epsilon$ for all $j\neq i$ when modifying region $i$.
Result: $\epsilon=0$ (per-region pressure is separable).
Rationale: All three per-region components depend only on the region’s own content:
  1. 1.
    Modifying region $i$’s schedule changes which meetings occupy region $i$’s slots.
  2. 2.
    This affects region $i$’s gap_ratio, overlap_count, and utilization_variance.
  3. 3.
    It has zero effect on any other region $j$’s pressure, since $j$’s components depend only on meetings within $j$’s time slots.
Note on attendee constraints: While the same attendee may have meetings in multiple time blocks, our overlap sensor counts overlaps within each block independently. Moving a meeting from region $i$ to region $j$ may create or resolve overlaps in region $j$, but this is a local effect on $j$’s pressure, not a coupling effect where modifying region $i$ affects region $j$ through some indirect mechanism.
B.3 Empirical Validation
Analysis of 270 pressure-field trials confirms separability:
  • Total tick-to-tick transitions analyzed: 9,873
  • Transitions with pressure improvement: 1,160 (11.7%)
  • Transitions with pressure degradation: 0 (0.0%)
  • Transitions with no change: 8,713 (88.3%)
The complete absence of pressure degradation is consistent with separable pressure: each accepted patch reduces local pressure, which directly reduces global pressure without adverse effects on other regions. (Patches that would increase local pressure are rejected by validation.)
Mean improvement magnitude: 2.67 pressure units. For context, initial problem pressure averages 12.7 units (range 0–42), so each successful patch reduces pressure by approximately 21% of the typical starting value. This substantially exceeds any plausible coupling bound, confirming that local improvements reliably translate to global improvements.
Verification of $\delta_{\min}>0$: Theorem 5.1 requires $\delta_{\min}>(n-1)\epsilon$. With $\epsilon=0$ (separable pressure), this reduces to $\delta_{\min}>0$. Empirically, the minimum observed pressure reduction across all 1,160 accepted patches is 1.0 pressure units, confirming $\delta_{\min}\geq 1>0$. The theorem’s convergence guarantee therefore applies to this domain.
Appendix C Theorem Proofs
This appendix provides complete proofs for the theoretical results in Section 4. Theorem 5.2 (Basin Quality) is omitted here as its proof is trivial and appears in full in the main text.
C.1 Proof of Theorem 5.1 (Convergence)
Proof.
We show that the system reaches a stable basin within the stated bound.
Setup. Let $P^{t}=P(s^{t})$ denote global pressure at tick $t$. Under $\epsilon$-bounded coupling, when a patch reduces local pressure $P_{i}$ by $\delta_{i}$, global pressure changes by:
$P^{t+1}-P^{t}=-\delta_{i}+\sum_{j\neq i}(P_{j}(s^{t+1})-P_{j}(s^{t}))$
By bounded coupling, $|P_{j}(s^{t+1})-P_{j}(s^{t})|\leq\epsilon$ for each $j\neq i$. With $n$ regions:
$P^{t+1}-P^{t}\leq-\delta_{i}+(n-1)\epsilon$
Progress guarantee. If $\delta_{i}\geq\delta_{\min}$ and $\delta_{\min}>(n-1)\epsilon$, then:
$P^{t+1}-P^{t}\leq-\delta_{\min}+(n-1)\epsilon<0$
Each tick with an applied patch reduces global pressure by at least $\delta_{\min}-(n-1)\epsilon>0$.
Termination. Since $P(s)\geq 0$ for all states and pressure decreases by a fixed minimum $\Delta=\delta_{\min}-(n-1)\epsilon$ per tick (when patches are applied), the system must reach a state where no region exceeds $\tau_{\text{act}}$ within:
$T\leq\frac{P_{0}}{\Delta}=\frac{P_{0}}{\delta_{\min}-(n-1)\epsilon}$
ticks. At this point, no region activates, and the system has reached a stable basin.
Stability maintenance. The decay constraint $\Delta_{f}>1-e^{-\lambda_{f}\cdot\tau_{\text{inh}}}$ ensures that fitness reinforcement from the final patches exceeds decay during the inhibition period. This prevents immediate re-activation after reaching stability. ∎
C.2 Proof of Theorem 5.3 (Basin Separation)
Proof.
We show that distinct stable basins are separated by pressure barriers of height at least $\tau_{\text{act}}$.
Basin definition. A stable basin $B$ is a connected region of state space where all states satisfy $P_{i}(s)<\tau_{\text{act}}$ for all regions $i$. Within a basin, no agent takes action (pressure below activation threshold).
Transition requirement. To move from basin $B_{1}$ to a distinct basin $B_{2}$, the system must pass through states not in either basin. At the boundary of $B_{1}$, at least one region $i$ must have $P_{i}(s)\geq\tau_{\text{act}}$ (otherwise, by continuity under separable pressure, the state would still be in $B_{1}$).
Barrier height. The minimum pressure exceedance required to exit a basin is $\tau_{\text{act}}$. This defines the barrier height separating basins.
Implication for decay. Without decay, once the system enters a basin, it remains there indefinitely—fitness remains high, preventing re-activation. Decay erodes fitness over time, eventually allowing pressure to exceed $\tau_{\text{act}}$ and enabling transition to potentially lower-pressure basins. ∎
C.3 Proof of Theorem 5.4 (Linear Scaling)
Proof.
We analyze the per-tick computational complexity.
Signal computation. For each of $m$ regions, computing $d$-dimensional signals requires $O(d)$ operations (reading local content, computing quality metrics). Total: $O(m\cdot d)$.
Pressure computation. For each region, computing pressure from $k$ signal axes requires $O(k)$ operations (weighted sum). Total: $O(m\cdot k)$.
Patch proposal. Each of $a$ actors proposes patches for activated regions. In the worst case (all regions activated), this is $O(m\cdot a)$ proposals. Each proposal involves an Large Language Model call, which dominates wall-clock time but is constant per call.
Selection. Sorting $O(m\cdot a)$ candidate patches by quality requires $O(m\cdot a\cdot\log(m\cdot a))$ comparisons. Greedy selection of non-conflicting patches is $O(m\cdot a)$ with appropriate data structures.
Coordination overhead. Critically, agents share no messages. Each agent reads shared artifact state (read-only) and writes proposed patches to a central queue. No inter-agent communication occurs. The fork pool for parallel validation has fixed size $K$, contributing $O(K)$ overhead independent of agent count.
Total. Summing components: $O(m\cdot d+m\cdot k+m\cdot a+m\cdot a\cdot\log(ma)+K)=O(m\cdot(d+k+a\cdot\log(ma)))$, independent of agent count $n$ beyond its contribution to $a$. ∎
C.4 Proof of Theorem 5.5 (Parallel Convergence)
Proof.
We extend Theorem 5.1 to the parallel setting.
Setup. Suppose $K$ patches are validated in parallel per tick, targeting disjoint regions $i_{1},\ldots,i_{K}$.
Combined pressure reduction. Under $\epsilon$-bounded coupling, each patch $k$ targeting region $i_{k}$ reduces local pressure by at least $\delta_{\min}$. The global pressure change from patch $k$ is:
$\Delta P_{k}\leq-\delta_{\min}+(n-1)\epsilon$
When patches target disjoint regions, their effects on global pressure are additive (no double-counting of coupling effects within the modified regions). The combined reduction is:
$P^{t+1}-P^{t}\leq\sum_{k=1}^{K}\Delta P_{k}\leq K\cdot(-\delta_{\min}+(n-1)\epsilon)$
Convergence bound. With combined reduction $K\cdot(\delta_{\min}-(n-1)\epsilon)$ per tick:
$T\leq\frac{P_{0}}{K\cdot(\delta_{\min}-(n-1)\epsilon)}$
Conflict handling. When multiple patches target the same region, only one is selected (highest quality). This reduces effective parallelism but maintains correctness—the selected patch still provides at least $\delta_{\min}-(n-1)\epsilon$ reduction. ∎